OpenStudy (anonymous):

What is the solution to the system of equations below: 2x+y-z=5 x+3z=14 -2x-3y+2z=2

OpenStudy (anonymous):

u givin me a medal?

OpenStudy (anonymous):

um sure. why not.

OpenStudy (mathmale):

Well, for starters, you have here a system of three linear equations in three unknowns, which implies that you have enough information to solve for the values of x, y and z.

OpenStudy (mathmale):

Which method(s) of solution is/are most familiar to you? Addition/subtraction? Substitution? Cramer's Rule? Matrices?

OpenStudy (anonymous):

addition/subtraction i do believe

OpenStudy (mathmale):

OK, Princess. Suppose you add the first and third equations together, term by term. The x terms will disappear. Willing to try that?

OpenStudy (anonymous):

sure! :D (and yes I am actually a real princess xD)

OpenStudy (mathmale):

Is your crown of gold or platinum?

OpenStudy (anonymous):

z=3, y=-2, x=5

OpenStudy (anonymous):

hheres the workings

OpenStudy (anonymous):

silver actually with lots of diamonds. When I added them together I got -2y+z=7

OpenStudy (mathmale):

Trolls will rule better if trolls will let Princess do her own work. Princess, that first result of yours is perfect.

OpenStudy (anonymous):

hahaha I like u mathmale xD Awesome so what's my next step?

OpenStudy (mathmale):

Now, Pricess Silver, please multiply the 2nd equation by 2 and add the resulting equation to the third equation. Again, the x terms will disappear.

OpenStudy (anonymous):



OpenStudy (mathmale):

Wouldn't it be more helpful to Princess, Troll, if you were to guide her to finding the solution herself?

OpenStudy (anonymous):

no, that is just too confusing

OpenStudy (anonymous):

just a second

OpenStudy (anonymous):

when she want to revise, she will be scrolling up and down

OpenStudy (mathmale):

Every Troll is entitled to his/her own opinion. Of course, Princess!

OpenStudy (anonymous):

definitely, just like gandhi

OpenStudy (anonymous):

but i am mandela while u are obama, we are both math lovin but different in terms of apporach

OpenStudy (mathmale):

So, Princess, you've already obtained -2y + 1z = 7 and are well on your way to obtaining a second equation in y and z only. Please type that equation here when you're finished.

OpenStudy (anonymous):

actually, can u generalise this system u talking abt? or do u wanna hear my generalization?

OpenStudy (mathmale):

I eagerly await Princess' response.

OpenStudy (anonymous):

sorry my mom was talking to me :P So, mathmale, I did what you said (since i had an explanation sorry trollswillrule but I need this for my final :D) and this is what I got: (x+3z=14)2= 2x+6z=28 then I added that to -2x-3y+2x=2 and got -3y+8z=30

OpenStudy (mathmale):

Perfect! So now you have two equations in two unknows: -2y + 1z = 7 -3y + 8z =30 I'd bet you know how to solve this simpler system.

OpenStudy (anonymous):

we would just have to times the first equation by 3 and the second equation by 2 to eliminate y correct?

OpenStudy (anonymous):

wait! Times the first equation by -3 so they'll cancel!

OpenStudy (mathmale):

Sounds like a great plan, Princess, except use -3 and 2 or 3 and -2. I think the reason for this will become apparent as you go along.

OpenStudy (mathmale):

Cool. Great minds think alike.

OpenStudy (anonymous):

so z=3... SWEETNESS! I can plug that into x+3z=14 and get x! :D:D:D:D

OpenStudy (mathmale):

How cool. Right you are! Be certain to check your solution (x, -2, 3) into one or more of the original equations to verify that it actually satisfies those equations.

OpenStudy (anonymous):

x=5 doesn't it? and they work! We are golden! Yes I might just pass this class after all :P

OpenStudy (mathmale):

Good luck on that exam tomorrow! Invite me to your coronation.

OpenStudy (anonymous):

Why thank you! And will do ;) Queen Holli might not be to excited that I'm planning on taking her place as Queen though xD

OpenStudy (mathmale):

I haven't actually found x myself. More important to me to help you solve the problem. Too bad for Q. Holli.

OpenStudy (anonymous):

gotcha :D Truth! lol thanks so much!

OpenStudy (mathmale):

My great pleasure. By all means post further questions and perhaps we'll collaborate again. Good night!